graph the function
f(x)=-1/2|x-4|+5

Answer by drk(1908)

Since this is absolute value function, we should be able to tell 3 things just by looking at the function:
(i) vertex
(ii) open up or down
(iii) wide, regular, skinny graph
---
(i) The vertex is at (4,5)
(ii) It opens down, because there is a (-) sign in front of the absolute value.
(iii) It is wide since there is a 1/2 in front of the absolute value.
--
to graph, plot (4,5). Then pick x = 2. We get
-1/2|2-4|+5 = 4. So, (2.4)
Now pick x = 6.
-1/2|6-4|+5 = 4. So, (6,4).
Make a rough sketch and you are done.